# Multi-Rank Subspace Change-Point Detection for Monitoring Robotic Swarms

### v1

@@ Abstract
We study the problem of real-time detection of covariance structure changes in high-dimensional streaming data, motivated by applications such as robotic swarm monitoring. Building upon the spiked covariance model, we propose the multi-rank Subspace-CUSUM procedure, which extends the classical CUSUM framework by tracking the top principal components to approximate a likelihood ratio. We provide a theoretical analysis of the proposed method by characterizing the expected detection statistics under both pre- and post-change regimes and offer principled guidance for selecting the drift and threshold parameters to control the false alarm rate. The effectiveness of our method is demonstrated through simulations and a real-world application to robotic swarm behavior data.
@@ 1 Introduction
The spiked covariance model (johnstone2001distribution, ) provides a tractable framework for capturing low-dimensional changes in covariance. It assumes that most of the variability is concentrated in a small number of principal components, making it particularly well suited to characterizing emerging signal subspaces. In this model, the covariance matrix consists of an identity baseline plus a low-rank component capturing the dominant subspace. Prior works have developed sequential detectors for rank-1 spike changes, exploiting either the largest sample eigenvalue (xie2020sequential, ) or an estimated leading eigenvector to form CUSUM-type procedures (xie2018first, ). However, real-world changes often involve multiple concurrent spikes of moderate strength, motivating a generalization to the multi–rank case. To address this, we extend the change-point detection framework by accommodating a general spiked covariance setting with an unknown multi-dimensional signal subspace.
@@
  • A multi–rank Subspace-CUSUM algorithm that tracks the top $d$ principal directions in a rolling window and raises an alarm when their combined projection energy exceeds a threshold.
  • Theoretical analyses showing how to set the drift parameter and detection threshold to achieve a prescribed average run length.
  • Numerical validations on both synthetic and real-world datasets, demonstrating that the proposed method effectively detects multi-spike changes and performs robustly even when individual signals are weak.
1.1 Related works
@@
Early approaches to change-point detection treated covariance shifts within the classical statistical process control. Traditional Shewhart control charts and CUSUM procedures (page1954continuous, ) were extended to monitor covariance changes under low-dimensional settings. For instance, Hotelling’s $T^{2}$ control chart (hotelling1947multivariate, ) can detect general mean or covariance shifts, and methods based on determinants or generalized likelihood have been used to detect known covariance changes. Also, healy1987note studied a multivariate CUSUM procedure for a covariance shift scaled by a constant. Similarly, edward1985multivariate proposed using the determinant of the sample covariance as a univariate statistic to flag a covariance change. However, these classical techniques often assumed either that the post-change covariance was fully specified or that data dimensions were modest.
@@
More recent works have tackled covariance change detection in high-dimensional contexts with unknown change-points in an offline manner. Notably, chen2004statistical developed hypothesis tests for a change in a Gaussian covariance, using information criteria to estimate the change-point. wang2021optimal proposed a binary segmentation approach to localize multiple covariance breaks, while avanesov2018change introduced a test based on differences in inverse covariance matrices. arias2012detection considered likelihood ratio tests for the appearance of a correlation between a specific subset of variables.
@@
Despite these advances in offline analysis, the problem of sequentially detecting covariance structure changes in real time remains relatively underexplored. jiao2018subspace studied the sequential detection of a subspace switch using a CUSUM-type procedure; however, their method requires knowing the post-change subspace in advance. Recently, dornemann2024detecting took a random matrix theory perspective to devise a hypothesis test for changes in the leading eigenvalues of a high-dimensional covariance sequence. zhang2023spectral Our work builds most directly on xie2018first ; xie2020sequential , who introduced sequential procedures for detecting when the covariance matrix shifts to a spiked model with an unknown principal component. In particular, they proposed a sliding-window largest eigenvalue chart and a subspace CUSUM detector on a single spike model.
@@
Moreover, many methods have been developed for detecting structural changes in evolving networks. zhang2023spectral proposed a spectral-CUSUM method to monitor changes in community structure by inspecting the eigen-decomposition of graph-related matrices. peel2015detecting introduced a technique to flag shifts in the large-scale connectivity pattern of an evolving network, and wang2017fast proposed a multi-step algorithm for detection in dynamic social networks. marangoni2014sequential developed a sequential approach based on a spectral CUSUM statistic for detecting the emergence of communities in networks. sulem2024graph designed a data-driven Siamese graph neural network, enabling online detection of distributional changes in a network without prior model knowledge. wang2023multilayer proposed an online tensor-based detection algorithm to identify changes in the latent connectivity structure across layers. Notably, chen2023discovering introduced a spectral iso-mirror embedding to reduce a dynamic network series to a scalar representation, and chen2024euclidean demonstrated that a spectral Euclidean mirror technique can effectively localize “first-order” change-points in network evolution.
@@ 2 Problem setup
In this section, we formalize two closely related formulations. In the emerging subspace setting, the data arrive as pure noise until an unknown time $\tau$ when a low-dimensional spike is added to the covariance. In the switching subspace problem, the nominal spike shifts to an unknown direction at time $\tau$. Both problems are naturally modeled by a spiked covariance model (johnstone2001distribution, ).
In the emerging subspace formulation, the observations follow a Gaussian distribution with identity covariance prior to the change, and a spiked covariance structure after the change:
$\begin{split}&x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}),\quad t=1,\dots,\tau,\\ &x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}+U\Lambda U^{\top}),\quad t=\tau+1,\tau+2,\dots,\end{split}$ (1)
where $\tau$ is the unknown change-point. The matrix $U\in\mathbb{R}^{k\times d}$ contains $d$ orthonormal columns representing the directions of the emerging signal subspace, and $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{d})$ is a diagonal matrix of signal strengths, with $\lambda_{1}\geq\cdots\geq\lambda_{d}>0$. The scalar $\sigma^{2}$ denotes the noise variance, which is considered known as it can be well-estimated by nominal data. We define the componentwise signal-to-noise ratio (SNR) as $\rho_{i}=\lambda_{i}/\sigma^{2}$ for $i=1,\dots,d$.
@@
$\begin{split}&x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}+U_{1}\Lambda U_{1}^{\top}),\quad t=1,\dots,\tau,\\ &x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}+U_{2}\Lambda U_{2}^{\top}),\quad t=\tau+1,\tau+2,\dots,\end{split}$ (2)
Here, $U_{1},U_{2}\in\mathbb{R}^{k\times d}$ are both semi-orthogonal matrices representing the pre- and post-change signal subspaces, satisfying $U_{i}^{\top}U_{i}=I_{d}$ for $i=1,2$. In this setting, the baseline subspace $U_{1}$ is assumed to be known a priori from the nominal distribution.
It is worth noting that the switching subspace model can be reduced to the emerging subspace formulation by projecting the data onto the orthogonal complement of $U_{1}$. Specifically, let $Q\in\mathbb{R}^{(k-d)\times k}$ be any semi-orthogonal matrix satisfying $QU_{1}=0$ and $QQ^{\top}=I_{k-d}$. Define the projected observations as
@@
Then the sequence ${y_{t}}$ is a random vector with distribution $N(0,\sigma^{2}I_{k-d})$ before the change, and $N(0,\sigma^{2}I_{k-d}+QU_{2}\Lambda U_{2}^{\top}Q^{\top})$ after the change. Denoting
$U=QU_{2}/\lVert QU_{2}\rVert_{F},\quad\tilde{\Lambda}=\Lambda\lVert QU_{2}\rVert_{F}^{2}=\left(d-\lVert U_{2}^{T}U_{1}\rVert^{2}_{F}\right)\Lambda,$
we recover an emerging subspace model of the form (LABEL:eq:emerging_subspace) for the projected observations as follows:
$\begin{split}&y_{t}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,\sigma^{2}I_{k-d}),\quad t=1,\dots,\tau,\\ &{y_{t}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,\sigma^{2}I_{k-d}+U\tilde{\Lambda}U^{T}),\quad t=\tau+1,\tau+2,\dots.}\end{split}$ (3)
@@
Thus, the switching subspace problem can be treated within the same analytical framework as the emerging subspace case. We note that this reduction involves a loss of information due to projection onto a lower-dimensional subspace. Nonetheless, it enables the use of a unified, computationally efficient detection strategy that applies to both emerging and switching subspace scenarios.
Additionally, the new signal power $\tilde{\Lambda}$ depends on the angles between the bases making up $U_{1}$ and $U_{2}$. In particular, there exist angles $20\leq\phi_{1}\leq\phi_{2}\leq\cdots\leq\phi_{d}\leq\pi/$ 2 such that the singular values of the $d\times d$ matrix $U_{2}^{T}U_{1}$ are $\cos\phi_{1},\cos\phi_{2},\dots,\cos\phi_{d}$. These $\phi_{i}$ are the principal angles between the two $d$-dimensional subspaces. That is,
$\tilde{\Lambda}\Lambda^{-1}=(d-\|U_{2}^{T}U_{1}\|_{F}^{2})I_{d}=\Big{(}d-\sum_{i=1}^{d}\cos^{2}\phi_{i}\Big{)}I_{d}=\Big{(}\sum_{i=1}^{d}\sin^{2}\phi_{i}\Big{)}I_{d},$
which quantifies the portion of the new subspace $U_{2}$ that lies outside the span of $U_{1}$.
@@ 3 Multi-rank Subspace-CUSUM
3.1 Exact CUSUM
The CUSUM test (page1954continuous, ) uses the log-likelihood ratio between the post-change and pre-change densities. The CUSUM statistic is defined as the maximum of the accumulated log-likelihood ratio over all possible change-point locations up to time $t$:
@@ 3.1 Exact CUSUM
where $f_{\infty}(\cdot)$ and $f_{0}(\cdot)$ denote the probability density functions of an observation before and after the change, respectively. Equivalently, $S_{t}$ can be computed recursively as (moustakides1986optimal, )
@@
where $x^{+}=\max\{0,x\}$. That is, at each time we add the new log-likelihood ratio and reset to zero if the sum becomes negative. The stopping time for the CUSUM procedure is the first time $t$ that $S_{t}$ exceeds a preset threshold $b$ chosen to control the false alarm rate:
@@
Using the Woodbury matrix identity and properties of determinants, the instantaneous log-likelihood ratio for a new observation $x_{t}$ is obtained as
$\log\frac{f_{0}(x_{t})}{f_{\infty}(x_{t})}$ $=\log\left[\frac{\left[(2\pi)^{k}\absolutevalue{\sigma^{2}I_{k}+U\Lambda U^{T}}\right]^{-1/2}\exp\left\{-\frac{1}{2}x_{t}^{T}\left(\sigma^{2}I_{k}+U\Lambda U^{T}\right)^{-1}x_{t}\right\}}{\left[(2\pi)^{k}\sigma^{2k}\right]^{-1/2}\exp\left\{-\frac{1}{2\sigma^{2}}x_{t}^{T}x_{t}\right\}}\right]$
@@
This exact CUSUM is known to be the oracle procedure (moustakides1986optimal, ). Notably, the update term has negative drift before change and positive drift after change. Indeed, letting $\mathbb{E}_{\infty}[\cdot]$ denote expectation under the pre-change distribution and $\mathbb{E}_{0}[\cdot]$ the expectation under the post-change distribution, one can see that (by Shannon’s inequality)
$\mathbb{E}_{\infty}\log\left(\frac{f_{0}}{f_{\infty}}\right)=-\frac{1}{2}\sum_{i=1}^{d}\Big{[}\log(1+\rho_{i})-\frac{\rho_{i}}{1+\rho_{i}}\Big{]}<0,$
@@
$\mathbb{E}_{0}\log\left(\frac{f_{0}}{f_{\infty}}\right)=\frac{1}{2}\sum_{i=1}^{d}\left[\rho_{i}-\log(1+\rho_{i})\right]>0.$
@@
3.2 Multi-rank Subspace-CUSUM procedure

### v2

# Multi-Rank Subspace Change-Point Detection with Application in Monitoring Robotic Swarms

@@ Abstract
We study real-time detection of low-rank changes in the covariance structure of high-dimensional streaming data, motivated by robotic swarm monitoring. Building on the spiked covariance model, we propose the Multi-rank Subspace-CUSUM (MRS-C) procedure, which extends classical CUSUM by tracking projection energy onto an estimated signal subspace. We analyze performance by characterizing the expected detection delay (EDD) under a prescribed average run length (ARL), deriving closed-form asymptotically optimal choices of the window size and drift. We further prove that MRS-C is first-order asymptotically optimal relative to the oracle Exact CUSUM, with an explicit efficiency constant that depends on heterogeneity in spike strengths. When the signal rank is unknown, we use a parallel procedure. Simulations and robotic swarm-behavior data illustrate robustness and effectiveness.
@@ 1 Introduction
The spiked covariance model (johnstone2001distribution) provides a tractable framework for capturing low-dimensional changes in covariance. It assumes that most of the variability is concentrated in a small number of principal components, making it particularly well suited to characterizing emerging signal subspaces. In this model, the covariance matrix consists of an identity baseline plus a low-rank component capturing the dominant subspace. Prior work has developed sequential detectors for rank-1 spike changes, exploiting either the largest sample eigenvalue (xie2020sequential) or an estimated leading eigenvector to construct CUSUM-type procedures (xie2018first). However, real-world changes often involve multiple concurrent spikes of moderate strength, motivating a generalization to the multi-rank case. To address this, we extend the change-point detection framework to accommodate a general spiked-covariance setting with an unknown multidimensional signal subspace.
@@
  • A multi-rank Subspace-CUSUM algorithm that effectively detects covariance structure changes by tracking the projection energy of observations onto an estimated signal subspace.
  • Rigorous theoretical analysis and optimality proof, establishing that the proposed procedure achieves first-order asymptotic optimality relative to the oracle Exact CUSUM. We also explicitly quantify the efficiency-loss constant arising from heterogeneity in signal strengths.
  • Principled guidance for parameter selection, derived from the asymptotic analysis of the expected detection delay (EDD). We provide closed-form expressions for the asymptotically optimal sliding-window size and drift parameter, thereby minimizing detection delay for a target false-alarm rate.
  • Comprehensive validation on both synthetic and real-world datasets, demonstrating that the proposed method outperforms existing baselines in multi-spike scenarios and successfully detects formation changes in robotic swarm behavior.
Compared with prior work, the main extension from xie2020sequential to the current work is that we allow the rank of the signal subspace to exceed one, which is essential for applications like the swarm monitoring considered here. We follow their general subspace-tracking-based detection framework, but extend it to handle multiple significant eigenvalues through a parallel procedure. Theoretically, our results are generalized to cover arbitrary fixed ranks.
1.1 Related work
@@
Early approaches to change-point detection treated covariance shifts within the classical statistical process control. Traditional Shewhart control charts and CUSUM procedures (page1954continuous) were extended to monitor changes in covariance in low-dimensional settings. For instance, Hotelling’s $T^{2}$ control chart (hotelling1947multivariate) can detect general mean or covariance shifts, and methods based on determinants or generalized likelihood have been used to detect known covariance changes. Also, healy1987note studied a multivariate CUSUM procedure for a covariance shift scaled by a constant. Similarly, edward1985multivariate proposed using the determinant of the sample covariance as a univariate statistic to flag a covariance change. However, these classical techniques often assumed either that the post-change covariance was fully specified or that the number of data dimensions was modest.
@@
More recent works have addressed covariance change detection in high-dimensional settings with unknown change-points in an offline manner. Notably, chen2004statistical developed hypothesis tests for a change in a Gaussian covariance, using information criteria to estimate the change-point. wang2021optimal proposed a binary segmentation approach to localize multiple covariance breaks, while avanesov2018change introduced a test based on differences in inverse covariance matrices. arias2012detection considered likelihood ratio tests for the appearance of a correlation between a specific subset of variables.
@@
Despite these advances in offline analysis, the problem of sequentially detecting changes in covariance structure in real time remains relatively underexplored. jiao2018subspace studied the sequential detection of a subspace switch using a CUSUM-type procedure; however, their method requires knowing the post-change subspace in advance. Recently, dornemann2026detecting took a random matrix theory perspective to devise a hypothesis test for changes in the leading eigenvalues of a high-dimensional covariance sequence. Our work builds most directly on xie2018first; xie2020sequential, who introduced sequential procedures for detecting when the covariance matrix shifts to a spiked model with an unknown principal component. In particular, they proposed a sliding-window chart of the largest eigenvalue and a subspace-tracking CUSUM-type detector for a single-spike model.
@@
Moreover, numerous methods have been developed to detect structural changes in evolving networks. zhang2023spectral proposed a spectral CUSUM method to monitor changes in community structure by inspecting the eigen-decomposition of graph-related matrices. peel2015detecting introduced a technique to flag shifts in the large-scale connectivity patterns of an evolving network, and wang2017fast proposed a multi-step algorithm for detecting such shifts in dynamic social networks. marangoni2014sequential developed a sequential approach based on a spectral CUSUM statistic for detecting the emergence of communities in networks. sulem2024graph designed a data-driven Siamese graph neural network, enabling online detection of distributional changes in a network without prior model knowledge. wang2026multilayer proposed an online tensor-based detection algorithm to identify changes in the latent connectivity structure across layers. Notably, chen2023discovering introduced a spectral iso-mirror embedding to reduce a dynamic network series to a scalar representation, and chen2024euclidean demonstrated that a spectral Euclidean mirror technique can effectively localize “first-order” change-points in network evolution.
Our work is closely related to recent modeling efforts on swarm dynamics, which provide a foundation for the datasets considered here. In theoretical swarm dynamics, introducing range-dependent communication delays has been shown to induce rich bifurcation structures, such as torus bifurcations that transform periodic rotational swarming into quasi-periodic collective motion (schwartz2020torus). Likewise, colliding swarms exhibit a critical coupling threshold beyond which they form a single coherent milling state instead of scattering, corresponding to a saddle-node bifurcation in the swarm-on-swarm dynamics (hindes2021critical). Extending these foundations, kamimoto2023chaotic demonstrated that when two swarms collide and merge into a combined milling state, their collective motion can transition from periodic to chaotic as key interaction parameters are varied.
@@ 2 Problem setup
In this section, we formalize two closely related formulations. In the emerging subspace setting, the data arrive as pure noise until an unknown time $\tau$ when a low-dimensional spike is added to the covariance. In the switching subspace problem, the nominal spike shifts to an unknown direction at time $\tau$. Both problems are naturally modeled by a spiked covariance model (johnstone2001distribution).
In the emerging subspace formulation, the observations $x_{t}\in\mathbb{R}^{k}$ follow a Gaussian distribution with identity covariance prior to the change, and a spiked covariance structure after the change:
$\begin{split}&x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}),\quad t=1,\dots,\tau,\\ &x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}+U\Lambda U^{T}),\quad t=\tau+1,\tau+2,\dots,\end{split}$ (1)
where $\tau$ is the unknown change-point. The matrix $U\in\mathbb{R}^{k\times d}$ contains $d$ orthonormal columns representing the directions of the emerging signal subspace, and $\Lambda=\operatorname{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{R}^{d\times d}$ is a diagonal matrix of signal strengths, with $\lambda_{1}\geq\cdots\geq\lambda_{d}>0$. The scalar $\sigma^{2}>0$ denotes the noise variance, which is assumed known, as it can be well estimated from nominal data. We define the componentwise signal-to-noise ratio (SNR) as $\rho_{i}=\lambda_{i}/\sigma^{2}$ for $i=1,\dots,d$.
@@
$\begin{split}&x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}+U_{1}\Lambda U_{1}^{T}),\quad t=1,\dots,\tau,\\ &x_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2}I_{k}+U_{2}\Lambda U_{2}^{T}),\quad t=\tau+1,\tau+2,\dots,\end{split}$ (2)
Here, $U_{1},U_{2}\in\mathbb{R}^{k\times d}$ are both semi-orthogonal matrices representing the pre- and post-change signal subspaces, satisfying $U_{1}^{T}U_{1}=U_{2}^{T}U_{2}=I_{d}$. In this setting, the baseline subspace $U_{1}$ is assumed to be known a priori from the nominal distribution.
It is worth noting that the switching subspace model can be reduced to the emerging subspace formulation by projecting the data onto the orthogonal complement of $U_{1}$. Specifically, we can take a matrix $Q\in\mathbb{R}^{(k-d)\times k}$ that satisfies $QU_{1}=0$ and $QQ^{T}=I_{k-d}$, and define the projected observations
@@
Then, each $y_{t}$ is Gaussian with pre-change covariance $\sigma^{2}I_{k-d}$ and post-change covariance $\sigma^{2}I_{k-d}+(QU_{2})\Lambda(QU_{2})^{T}.$ Here, the post-change covariance matrix has rank at most $d$ and admits an eigendecomposition
$(QU_{2})\Lambda(QU_{2})^{\top}=U_{y}\Lambda_{y}U_{y}^{\top},$
where $U_{y}\in\mathbb{R}^{(k-d)\times r}$ has orthonormal columns, $\Lambda_{y}=\operatorname{diag}(\lambda_{y,1},\ldots,\lambda_{y,r})$ with $\lambda_{y,j}>0$, and $r\leq d$ is the effective post-change rank. In the case where $\operatorname{col}(U_{2})\cap\operatorname{col}(U_{1})=\{0\}$, we have $r=d$; otherwise, components of $U_{2}$ lying in $\operatorname{col}(U_{1})$ are annihilated by the projection, and the effective rank is reduced.
Therefore, after projection, the post-change distribution can always be written in standard spiked-covariance form as
$y_{t}\sim N(0,\sigma^{2}I_{k-d}+U_{y}\Lambda_{y}U_{y}^{T}),$
@@
with signal subspace $\operatorname{col}(U_{y})$ and spike strengths $\lambda_{y,1},\ldots,\lambda_{y,r}$. Thus, the switching subspace problem can be treated within the same analytical framework as the emerging subspace case. We note that this reduction involves a loss of information due to projection onto a lower-dimensional subspace. Nonetheless, it enables a unified, computationally efficient detection strategy applicable to both emerging and switching subspace scenarios. Accordingly, throughout the rest of this paper, we restrict attention to the emerging subspace model and develop the detection procedures and theoretical analysis in that setting, noting that the switching subspace problem can be handled via the projection-based reduction described above.
Remark 2.1 (Geometric interpretation via principal angles).
Let $\theta_{1},\ldots,\theta_{d}\in[0,\pi/2]$ denote the principal angles between the subspaces $\operatorname{col}(U_{1})$ and $\operatorname{col}(U_{2})$. Then the matrix
$U_{2}^{\top}(I-U_{1}U_{1}^{\top})U_{2}$
has eigenvalues $\sin^{2}(\theta_{1}),\ldots,\sin^{2}(\theta_{d})$. In particular, when $\Lambda=\lambda I_{d}$, the nonzero eigenvalues of the projected spike term $(QU_{2})\Lambda(QU_{2})^{\top}$ are given by
$\lambda_{y,j}=\lambda\sin^{2}(\theta_{j}),\qquad j=1,\ldots,d,$
with those equal to zero corresponding to $\theta_{j}=0$, i.e., directions in $\operatorname{col}(U_{2})$ that lie in $\operatorname{col}(U_{1})$ and are annihilated by the projection. More generally, for diagonal $\Lambda=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{d})$, the projected spike strengths are the eigenvalues of
$\Lambda^{1/2}\,U_{2}^{\top}(I-U_{1}U_{1}^{\top})\,U_{2}\,\Lambda^{1/2},$
which couples the original spike magnitudes with the relative orientation of the two subspaces.
2.1 Motivating application: robotic swarm monitoring
The emerging and switching subspace problems are directly relevant to modeling and analyzing collective dynamics in swarms. For example, consider a scenario in which a group of autonomous agents moves cohesively, forming stable formations characterized by low-dimensional subspaces in their positional and velocity data. When the swarm undergoes structural transitions such as splitting into subgroups, merging, or forming new coherent structures, the underlying covariance structure of the agents’ movements changes accordingly, resulting in the emergence or shift of principal subspaces.
The rationale for modeling structural changes via low-rank covariance shifts is grounded in the intrinsic dimensionality commonly observed in real-world data. In many high-dimensional applications, including image sequences, network behaviors, and swarm trajectories, temporal dynamics typically concentrate energy in a few dominant directions, with the remaining dimensions largely capturing noise or insignificant variation. As illustrated in (xie2016sequential, Figure 2), practical datasets often exhibit sharp transitions in spectral characteristics, where a small number of principal components abruptly emerge or vanish at a change-point. Leveraging a low-rank representation not only captures these fundamental structural patterns but also enhances sensitivity to subtle yet meaningful variations, thereby enabling quicker and more reliable change-point detection.
@@ 3 Multi-rank Subspace-CUSUM
Before presenting our Multi-rank Subspace-CUSUM (MRS-C), we briefly recall the classical CUSUM test proposed by page1954continuous. When the pre-change and post-change densities are both known, this Exact CUSUM test is optimal for sequential change detection (moustakides1986optimal). In our setting, this would require full knowledge of the post-change signal subspace $U$ and the spike magnitudes $\lambda_{i}$, quantities that are unavailable in practice. The Exact CUSUM is therefore an oracle benchmark. The MRS-C discussed in Section 3.2 estimates $U$ and $\lambda_{i}$ on the fly, enabling a practical procedure that approximates this oracle test.
3.1 Oracle benchmark: Exact CUSUM
The (Exact) CUSUM test (page1954continuous) uses the log-likelihood ratio between the post-change and pre-change densities. The CUSUM statistic is defined as the maximum of the accumulated log-likelihood ratio over all possible change-point locations up to time $t$:
@@ 3.1 Oracle benchmark: Exact CUSUM
where $f_{\infty}(\cdot)$ and $f_{0}(\cdot)$ denote the probability density functions of an observation before and after the change, respectively. Equivalently, $S_{t}$ can be computed recursively as (moustakides1986optimal)
@@
where $x^{+}=\max\{0,x\}$. That is, at each time we add the new log-likelihood ratio and reset it to zero if the sum becomes negative. The stopping time for the CUSUM procedure is the first time $t$ that $S_{t}$ exceeds a preset threshold $b$ chosen to control the false alarm rate:
@@
Under the emerging subspace model (LABEL:eq:emerging_subspace) that we study in this paper, the log-likelihood ratio for an observation $x_{t}$ is obtained as
$\log\frac{f_{0}(x_{t})}{f_{\infty}(x_{t})}$ $=\log\left[\frac{\left[(2\pi)^{k}\left|\sigma^{2}I_{k}+U\Lambda U^{T}\right|\right]^{-1/2}\exp\left\{-\frac{1}{2}x_{t}^{T}\left(\sigma^{2}I_{k}+U\Lambda U^{T}\right)^{-1}x_{t}\right\}}{\left[(2\pi)^{k}\sigma^{2k}\right]^{-1/2}\exp\left\{-\frac{1}{2\sigma^{2}}x_{t}^{T}x_{t}\right\}}\right]$
@@
Notably, the update term exhibits negative drift before the change and positive drift after the change. Indeed, letting $\mathbb{E}_{\infty}[\cdot]$ denote expectation under the pre-change distribution and $\mathbb{E}_{0}[\cdot]$ the expectation under the post-change distribution, one can see that
$\mathbb{E}_{\infty}\left[\log\frac{f_{0}(x_{t})}{f_{\infty}(x_{t})}\right]=-\frac{1}{2}\sum_{i=1}^{d}\Big[\log(1+\rho_{i})-\frac{\rho_{i}}{1+\rho_{i}}\Big]<0,$ (4)
@@
$\mathbb{E}_{0}\left[\log\frac{f_{0}(x_{t})}{f_{\infty}(x_{t})}\right]=\frac{1}{2}\sum_{i=1}^{d}\left[\rho_{i}-\log(1+\rho_{i})\right]>0.$ (5)
@@
3.2 The MRS-C procedure